# Prediction intervals for overdispersed binomial endpoints and their application to historical control data

### v2

@@ 5.2 Coverage probabilities for the simulation setting inspired by long term carcinogenicity studies
This part of the simulation was inspired by the estimates for the binomial proportion and overdispersion reported by Menssen and Schaarschmidt 2019 for HCD obtained from long term carcinogenicity (LTC) studies run with B6C3F1 mice. In order to assess the coverage probabilities for each of the six methods on a broad range of realistic parameter combinations, the simulation was run for all 96 combinations of $H={5,10,20,100}$, $\pi={0.01,0.1,0.2,0.3,0.4,0.5}$, $\phi={1.001,1.5,3,5}$ and a fixed cluster size of $n_{h}=n^{*}=50\quad\forall\quad h=1,\ldots,H$.
Due to the relatively small cluster size, the underlying distribution is heavily discrete and can become heavily right skewed for small proportions and/or high overdispersion. Hence, it is possible that in some settings (especially $\pi=0.01$) the data contains many zeros. If it happened that a sampled ”historical” data set contained only zeros ($y_{h}=0\quad\forall\quad h=1,\ldots H$), the first observation was set to $y_{1}=0.5$ and the corresponding cluster size was set to $n_{1}=49.5$ following Menssen and Schaarschmidt 2019. This procedure was applied to all frequentist methods (heuristics and prediction intervals) in order to enable the estimation process in this extreme settings, whereas otherwise, the estimates for the binomial proportion and overdispersion would become zero.
Bayesian hierarchical modeling did not lead to convergence regularly enough for reliable results to be reported. In particular, the precision hyperparameter related to overdispersion is very difficult to estimate, which also depends on the prior choice. Therefore, the GLMM is applied here as an alternative.
The heuristical methods (Fig. 3) behaved in a similar fashion than in the MNT-setting: With rising amount of historical control groups, the coverage probability of the range approaches one (Fig. 3A). In the absence of overdispersion, the np-Chart (Fig. 3B) yields coverage probabilities close to the nominal level, but with a rising amount of overdispersion, its coverage probabilities decrease below 0.6.
@@
Figure 3: Simulated coverage probabilities of heuristical methods (LTC-setting) A: Historical range, B: np-chart, C: Mean $\pm$ 2 SD, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$. Note the different scales of the y-axis of the sub-graphics (0.5-1 for A and B, 0.8-1 for C).
@@
With increasing overdispersion $\phi$ and deceasing proportions $\pi\leq 0.1$ the amount of zeros and hence the right skeweness of the underlying distribution increases. This means that the lower borders of all three prediction intervals tend to cover more future observations than desired. This is the reason why all three prediction intervals practically yield 97.5 % upper prediction borders for $\pi=0.01$ and the conservativeness of the lower borders increases for $\pi=0.1$ with an increase of $\phi$. This is not a misbehavior of the methodologies, but rather reflects a core feature of the underlying distribution.
@@ 7 Discussion
The validation of a concurrent control group based on prediction intervals is recommended by several authors (Menssen 2023, Kluxen et al. 2021) and since shortly, seems to be preferred by the European Food Safety Agency (EFSA 2024) over other (heuristical) methods. Nevertheless, since literature for the calculation of prediction intervals for clustered dichotomous data is relatively scarce, the practical application of HCL tends to rely on several heuristics, such as the application of the mean $\pm$ 2 SD (Rotolo et al. 2021, Prato et al. 2023, de Kort et al. 2020, Kluxen et al. 2021) or limits from Sheward control charts (Lovell et al. 2018, Dertinger et al. 2023). Our simulations have shown that these heuristical methods do not control the statistical error to a satisfactory level and hence, these methods can not be recommended for the practical validation of a current control group.
Valverde-Garcia et al. 2018 report HCD for avian reproduction and calculated prediction intervals for several dichotomous endpoints such as the number of viable vs. nonviable hatchlings. Anyhow, the applied prediction intervals depend on the assumption of univariate normality and hence do not match with the dichotomy of the assessed endpoints.
@@
Several prediction intervals are available from the literature that are based on the assumption of independent binomial distributed observations (Meeker et al. 2017), but ignore the clustured structure of the HCD (and hence the possibility for between-study overdispersion). Menssen and Schaarschmidt 2019 showed, that these intervals fail to approach the nominal coverage probability in the presence of overdispersion.
The magnitude of overdispersion reflects the between-study variation which is commonly present in toxicological HCD and hence, can be found in several published HC data sets in the literature (Tarone 1982a, Tarone 1982b, Levy et al. 2019, Menssen et al. 2024, Carlus et a.l 2013, Dertinger et al. 2023, Tug et al. 2024, see supplementary materials section 4). Therefore prediction intervals that are based on the simple binomial assumption can not be recommended for their application to HCD with a toxicological background. The need for methodology that account for between-study overdispersion was addressed by the proposed prediction intervals. All four methods explicitly account for the clustered structure of toxicological HCD and hence also for the possible presence of overdispersion.
@@
Both of the calibrated prediction intervals are improved versions of the quasi-binomial prediction interval of Menssen and Schaarschmidt 2019 which rely on a bootstrap procedure that adapts the formerly symmetrical prediction interval to possible skeweness of the underlying distribution. Note, that this bootstrap calibration procedure has already been applied successfully for the calculation of prediction intervals drawn from random effects models (Menssen and Schaarschmidt 2022) or for overdispersed count data (Menssen et al. 2024). This means that, as long as one is able to bootstrap new observations from a model and the prediction standard error $\widehat{se}(\hat{y}^{*}-Y^{*})=\sqrt{\widehat{var}(\hat{y}^{*})+\widehat{var}(Y^{*})}$ can be formulated, the presented bootstrap calibration procedure provides a standardized framework for the calculation of prediction intervals for various distributional assumptions and/or models. An implementation of the bisection used in step 8 and 9 of the bootstrap calibration is provided via the function bisection() of the predint R package.
@@
The prediction intervals drawn from Bayesian hierarchical models yield coverage probabilities that in the MNT-setting mainly remain below the nominal level (if overdispersion is present), but the model did not converge to most of the simulated data sets that mimic HCD from carcinogenicity studies (mainly if many zeros were present in the data). It is noteworthy, that in the Monte-Carlo simulations, Bayesian hierarchical models were fit, always using the same prior $\kappa\sim\text{Ga}(2,\>5\cdot 10^{-5})$ regardless of the simulation setting. But, from the application to the real life example it is obvious, that this type of model needs a case to case adaption of the prior distribution to provide a reasonable domain for the hyperparameter $\kappa$ (which depends on the amount of overdispersion). This might, at least partly, explain the liberality of this method. Unfortunately, the necessity for a case by case adaption of the prior distribution is far beyond the scope of most practitioners who are in charge of the reporting of HCD (mainly toxicologists that are barely trained in statistics). Hence this necessity might be a drawback for practical application.
@@
Contrary to the Bayesian hierarchical models, the Bayesian GLMM did not have convergence problems in the LTC-setting. But, the applied GLMM mainly yield prediction intervals with coverage probabilities above the nominal level (except for a higher amount of overdispersion). This is, because such models a priori assume overdispersion (or, equivalently, additional non-zero between-study variance) to be present, and estimate the respective parameters, even if they are actually not part of the data-generating process.
@@
An alternative to a fully Bayesian approach might be the application of an empirical Bayes procedure in which the hyperparameter’s domain is pre-estimated from the HCD, similar to the approaches of Tarone 1982a who incorporated HCD in a trend test or Kitsche et al. 2012 who exploit HCD to be accounted for in a Dunnett-type mean comparison. Anyhow, to the authors knowledge, an empirical Bayes procedure for the calculation of prediction intervals for overdispersed binomial data does not exist and might be a subject for future research.
@@
It is noteworthy, that the coverage probability of the calibrated intervals is mainly effected by the amount of available historical information and that in most of the cases the amount of overdispersion or the magnitude of the binomial proportion is of minor importance. This means, that for a high cluster size (as in the MNT-setting) both calibrated prediction intervals yield coverage probabilities satisfactorily close to the nominal level even for only five historical control groups. Contrary, for smaller cluster sizes (as in the LTC-setting) the number of historical control groups has to be slightly higher (say 10) to enable coverage probabilities close enough to the nominal level. Furthermore, note that with decreasing cluster size, the discreteness of the underlying distribution rises and hence, the nominal 0.95 coverage probability can not completely be achieved in all of the possible cases (Agresti 1998).
@@
Another relevant application is the evaluation of HCD with individual (non-aggregated) values reported (e.g., for six wells in one historical MNT-experiment), where extra-Binomial variability is also possible within one experiment. Here, GLMMs (both frequentist and Bayesian) as well as hierarchical models can be very helpful. Jeske and Young 2013 provided a bootstrap based procedure for the calculation of prediction intervals based on one-way GLMM, but to the authors knowledge, a general solution for the calculation of prediction intervals drawn from GLMM is not yet available. Hence, also this topic remains as an open issue.
@@ 8 Conclusions
  • Symmetrical HCL do not account for possible right- (or left-) skeweness of the data and hence do not ensure for equal tail probabilities.
@@
  • In most of the cases, the Bayesian prediction intervals do not yield coverage probabilities close enough to the nominal level (Bayesian hierarchical prediction intervals remain liberal, prediction intervals drawn from Bayesian GLMMM tend to behave conservatively).

### v3

@@ 5.2 Coverage probabilities for the simulation setting inspired by long term carcinogenicity studies
This part of the simulation was inspired by the estimates for the binomial proportion and overdispersion reported by Menssen and Schaarschmidt 2019 for HCD obtained from long term carcinogenicity (LTC) studies run with B6C3F1 mice. 96 combinations of values for the number of historical studies $H$, the binomial proportion $\pi$ and the amount of overdispersion $\phi$ were run, of which 54 combinations directly reflect the properties of the reported real life HCD (combinations of the bold faced parameter values in tab. 2), whereas the remaining 42 combinations of parameter values were run in order to enhance the generalisability of the simulation.
Table 2: Parameter values used for the LTC-inspired simulation
Parameter Values1
No. of historical studies ($H$) 5, 10, 20, 100
Binomial proportion ($\pi$) 0.01, 0.1, 0.2, 0.3, 0.4, 0.5
Overdispersion ($\phi$)2 1.001, 1.5, 3, 5
Historical cluster size ($n_{h}$) 50
Future cluster size ($n^{*}$) 50
Bold numbers: Parameter values that reflect estimates obtained from real life HCD. 1: Simulations were run for all 96 combinations of the given parameter values. 2: Since the cluster sizes $n_{h}$ and $n_{*}$ are constant with $\phi\widehat{=}1+(n_{h}-1)\rho\widehat{=}1+(n^{*}-1)\rho$, the given values of $\phi$ correspond to intra-class correlation coefficients $\rho$ of 2.04e-05, 0.01020, 0.04081, 0.08163.
Due to the relatively small cluster size, the underlying distribution is heavily discrete and can become heavily right skewed for small proportions and/or high overdispersion (see fig.7). Hence, it is possible that in some settings (especially $\pi=0.01$) the data contains many zeros. If it happened that a sampled ”historical” data set contained only zeros ($y_{h}=0\quad\forall\quad h=1,\ldots H$), the first observation was set to $y_{1}=0.5$ and the corresponding cluster size was set to $n_{1}=49.5$ following Menssen and Schaarschmidt 2019. This procedure was applied to all frequentist methods (heuristics and prediction intervals) in order to enable the estimation process in this extreme settings, whereas otherwise, the estimates for the binomial proportion and overdispersion would become zero.
Figure 5: Simulated coverage probabilities of heuristical methods (LTC-setting) A: Historical range, B: np-chart, C: Mean $\pm$ 2 SD, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$. Coverage probabilities below 0.6 are excluded from the graphic
For almost all simulation settings, Bayesian hierarchical modeling faced serious convergence problems. Therefore, the simulation results are heavily unreliable and not worth reporting. In particular, the precision hyperparameter related to overdispersion is very difficult to estimate, which also depends on the prior choice. Therefore, the only Bayesian method presented in this section is the prediction interval which is calculated based on a Bayesian GLMM.
The heuristical methods (Fig. 5) behaved in a similar fashion than in the MNT-setting: With rising amount of historical control groups, the future observation is always covered (Fig. 5A). In the absence of overdispersion, the np-Chart (Fig. 5B) yields coverage probabilities close to the nominal level, but with a rising amount of overdispersion, its coverage probabilities decrease below 0.6.
@@
With increasing overdispersion $\phi$ and deceasing proportions $\pi\leq 0.1$ the right skewness of the underlying distribution increases (see supplementary fig. 1 and 2) and its lower 2.5% quantile can approach zero (see fig. 7). With other words: in this case the lower prediction border aims to approximate a true quantile that is zero or close to zero.
This is the reason why all three prediction intervals practically yield 97.5 % upper prediction borders for $\pi=0.01$ and the conservativeness of the lower borders increases for $\pi=0.1$ with an increase of $\phi$. This is not a misbehavior of the methodologies, but rather reflects a core feature of the underlying distribution and one should think about the application of an upper prediction border, rather than a prediction interval if many zeros are present in the data.
Figure 7: Average limits vs. true quantiles of the underlying distribution (LTC-setting). Grey area: Probability mass function of the underlying beta binomial distribution. Black lines: True underlying 2.5 % and 97.5% quantiles. Dashed red lines: Average limits obtained from the simulation.
@@ 7 Discussion
The validation of a concurrent control group based on prediction intervals is recommended by several authors (Menssen 2023, Kluxen et al. 2021) and since shortly, seems to be preferred by the European Food Safety Agency (EFSA 2024) over other (heuristical) methods. Nevertheless, since literature for the calculation of prediction intervals for clustered dichotomous data is relatively scarce, the practical application of HCL tends to rely on several heuristics, such as the application of the mean $\pm$ 2 SD (Rotolo et al. 2021, Prato et al. 2023, de Kort et al. 2020, Kluxen et al. 2021) or limits from Shewhart control charts (Lovell et al. 2018, Dertinger et al. 2023). Our simulations have shown that heuristical methods do not control the statistical error to a satisfactory level and hence, can not be recommended for the practical validation of a current control group.
@@
Contrary to this, we could show, that prediction intervals satisfactorily control the statistical error, if calculated based on a model that match the statistical properties of the HCD. Several prediction intervals are available from the literature that are based on the assumption of independent binomial distributed observations (see e.g. Meeker et al. 2017 or Wang 2008), but ignore the clustered structure of the HCD (and hence the possibility for between-study overdispersion). In the context of HCD from long term carcinogenicity studies, Menssen and Schaarschmidt 2019 showed, that these intervals fail to approach the nominal coverage probability in the presence of overdispersion. This demonstrates that between-study overdispersion is a crucial part of the underlying data generating process and hence needs to be considered for HCL calculation.
The magnitude of overdispersion reflects the between-study variation which is commonly present in toxicological HCD and hence, can be found in several published HC data sets in the literature (Tarone 1982a, Tarone 1982b, Levy et al. 2019, Menssen et al. 2024, Carlus et a.l 2013, Dertinger et al. 2023, Tug et al. 2024, see supplementary materials section 4). However, in practice, systematic between-study variation might be induced by a mixture of both, controllable causes such as different handling of experimental units by different lab personnel, as well as uncontrollable causes such as the genetic variation inherent in the population of experimental units. This means that in practice, overdispersed HCD should be handled with care in a way that the magnitude of overdispersion reflects mainly the biological variability that is inherent of the biological process under observation, rather than the variability that is induced by controllable causes.
@@
Both of the calibrated prediction intervals are improved versions of the quasi-binomial prediction interval of Menssen and Schaarschmidt 2019 which rely on a bootstrap procedure that adapts the formerly symmetrical prediction interval to possible skewness of the underlying distribution. Note, that this bootstrap calibration procedure has already been applied successfully for the calculation of prediction intervals drawn from random effects models (Menssen and Schaarschmidt 2022) or for overdispersed count data (Menssen et al. 2024). This demonstrates that, as long as one is able to bootstrap new observations from a model and the prediction standard error $\sqrt{\widehat{var}(\hat{y}^{*})+\widehat{var}(Y^{*})}$ can be formulated, the presented bootstrap calibration procedure provides a standardized framework for the calculation of prediction intervals for various distributional assumptions and/or models. An implementation of the bisection used in step 8 and 9 of the bootstrap calibration is provided via the function bisection() of the R package predint.
@@
The prediction intervals drawn from Bayesian hierarchical models yield coverage probabilities that mainly remain below the nominal level in the MNT-setting (if overdispersion is present), but the model did not converge to most of the simulated data sets that mimic HCD from carcinogenicity studies (mainly if many zeros were present in the data). It is noteworthy, that in the Monte-Carlo simulations, Bayesian hierarchical models were fit, always using the same prior $\kappa\sim\text{Ga}(2,\>5\cdot 10^{-5})$ regardless of the simulation setting. But, from the application to the real life example it is obvious, that this type of model needs a case to case adaption of the prior distribution to provide a reasonable domain for the hyperparameter $\kappa$ (which depends on the amount of overdispersion). This might, at least partly, explain the liberality of this method. Unfortunately, the necessity for a case by case adaption of the prior distribution is far beyond the scope of most practitioners who are in charge of the reporting of HCD (mainly toxicologists that are barely trained in statistics). Hence this necessity might be a drawback for practical application.
@@
Contrary to the Bayesian hierarchical models, the Bayesian GLMM did not have convergence problems in the LTC-setting. But, the applied GLMM mainly yield prediction intervals with coverage probabilities above the nominal level (except for a higher amount of overdispersion). This is, because such models a priori assume overdispersion or, equivalently, additional non-zero between-study variability to be present, and hence provide estimates for the respective parameters, even if between-study variability is absent in the the data-generating process.
@@
An alternative to a fully Bayesian approach might be the application of an empirical Bayes procedure in which the hyperparameters domain is pre-estimated from the HCD, similar to the approaches of Tarone 1982a who incorporated HCD in a trend test or Kitsche et al. 2012 who exploit HCD to be accounted for in a Dunnett-type mean comparison. Anyhow, to the authors knowledge, an empirical Bayes procedure for the calculation of prediction intervals for overdispersed binomial data does not exist and might be a subject for future research.
It is noteworthy, that the coverage probability of the calibrated intervals is mainly effected by the amount of available historical information, defined by the numbers of experimental units within each control group $n_{h}$ and number of historical studies $H$. In most of the cases the amount of overdispersion or the magnitude of the binomial proportion is of minor importance (as long as one accepts that the lower border can become zero for low proportions and / or high overdispersion such that the desired 95 % prediction interval practically results in a 97.5 % upper prediction limit).
@@
This means, that for a high cluster size (as in the MNT-setting) both calibrated prediction intervals yield coverage probabilities satisfactorily close to the nominal level even for only five historical control groups. Contrary, for smaller cluster sizes (as in the LTC-setting) the number of historical control groups has to be slightly higher (say 10) to enable coverage probabilities close enough to the nominal level. Furthermore, note that with decreasing cluster size ($n_{h}$ and $n^{*}$), the discreteness of the underlying distribution rises and hence, the nominal 0.95 coverage probability can not completely be achieved in all of the possible cases (Agresti 1998).
@@
Another relevant application is the evaluation of HCD with individual (non-aggregated) values reported (e.g., for individual wells in one historical MNT-experiment rather than the complete plate), where extra-Binomial variability is also possible within one experiment (e.g. between the wells of the same plate). Here, GLMMs (both frequentist and Bayesian) as well as hierarchical models can be very helpful. Jeske and Young 2013 provided a bootstrap based procedure for the calculation of prediction intervals based on one-way GLMM, but to the authors knowledge, a general solution for the calculation of prediction intervals drawn from GLMM is not yet available. Hence, also this topic remains as an open issue that needs further investigation.
@@ 8 Conclusions
  • Symmetrical HCL do not account for possible right- (or left-) skewness of the data and hence do not ensure for equal tail probabilities.
@@
  • Due to their poor statistical properties (coverage probabilities below the nominal level) and the need for a case to case adaption of prior distributions, prediction intervals drawn from Bayesian hierarchical models can not be recommended for practical application.
  • Prediction intervals computed based on Bayesian GLMM have coverage probabilities which are relatively close to the nominal level (in most cases). Because they can be easily adapted to handle more than one level of hierarchy, they are promising candidates for the application to non-aggregated HCD.
@@
9 Acknowledgments
We have to thank Frank Schaarschmidt for fruitful discussions on prediction intervals and the impact of overdispersion on the inference drawn from dichotomous data. Furthermore, we have to thank both reviewers for their valuable comments which greatly helped to improve our manuscript.